### v2

# Twin electroweak bubble nucleation and gravitational wave under the $S_{3}$ symmetry of two-Higgs-doublet model

Abstract
Sphaleron electroweak phase transition (EWPT) is calculated in two phase transition stages, thereby showing that the twin (or double) bubble nucleation structure of the phase transition and gravitational wave is in the investigation area of future detectors. With $v^{2}=v^{2}_{1}+v^{2}_{2}$ ($v_{1}$ and $v_{2}$ are two vacuum average values (VEV)), the parameter $\tan\beta=v_{2}/v_{1}$, is the ratio between two VEVs although it does not affect the strength of EWPT but affect the sphaleron energy. However, it only causes this energy to increase slightly. As $a=v^{2}/v_{2}^{2}$ increases, the maximum difference of sphaleron energy in one stage is about $6.15$ TeV. $a$ affects the expansion of bubbles during two phase transitions. The more $a$ increases, the more the expansion of two bubbles is at the same time. This ratio does not greatly affect the sphaleron energy but has an impact on gravitational waves. The larger the masses of the charged Higgs particles are, the greater the gravitational wave energy density ($\Omega h^{2}$) is. When the frequency is in the range $0-1.2$ mHz, $\Omega h^{2}$ will has a maximum value in the range $10^{-12}-10^{-11}$ for all values of $a$ so this can be detected in the future.
pacs:
11.15.Ex, 12.60.Fr, 98.80.Cq
Keywords: Spontaneous breaking of gauge symmetries, Extensions of electroweak Higgs sector, Particle-theory models (Early Universe)
I INTRODUCTION
Particle physics today not only studies each individual problem in depth but also combines it with Cosmology and Astronomy. The baryon asymmetry is a problem of intersection among those fields. The possible causes of this problem are summarized and presented quite extensively. Indeed, for Higgs 125 GeV, Standard Model (SM) has only a second order electroweak phase transition mkn ; SME ; SMEb ; SMEc ; SMEd ; michela . Therefore, SM should be extended (for example, Refs. plv ; 2b ; 2c ; BSM ; BSMb ; majorana ; majoranab ; thdm ; thdmb ; ESMCO ; elptdm ; elptdma ; elptdmb ; elptdmc ; elptdmd ; phonglongvan ; phonglongvanb ; phonglongvan2 ; SMS ; dssm ; munusm ; lr ; singlet ; singletb ; singletc ; singletd ; mssm1 ; mssm1b ; mssm1c ; twostep ; twostepb ; twostepc ; 1101.4665 ; 1101.4665b ; 1101.4665c ; jjgb ; jjgc ; jjgd ; scaling1 ; scaling2 ; scaling3 ; Ahriche3 ; Ahriche3b ; Ahriche3c ; Ahriche3d ; Ahriche3e ; Fuyuto ; Fuyutob ; Fuyutoc ; span ; chr ; cde ; kusenko ). There are 3 main branches: the first is due to the existence of exotic particles (beyond the standard model (SM)); the second is because of parameters or interactions in SM changing at a electroweak transition temperature plv ; 2b ; 2c ; BSM ; BSMb ; majorana ; majoranab ; thdm ; thdmb ; ESMCO ; elptdm ; elptdma ; elptdmb ; elptdmc ; elptdmd ; phonglongvan ; phonglongvanb ; phonglongvan2 ; SMS ; dssm ; munusm ; lr ; singlet ; singletb ; singletc ; singletd ; mssm1 ; mssm1b ; mssm1c ; twostep ; twostepb ; twostepc ; 1101.4665 ; 1101.4665b ; 1101.4665c ; jjgb ; jjgc ; jjgd ; scaling1 ; scaling2 ; scaling3 ; Ahriche3b ; Ahriche3c ; Ahriche3d ; Ahriche3e ; Fuyutob ; Fuyutoc ; span ; chr ; cde ; kusenko ; the third be having more than 4 dimensions of spacetime pn .
Except for the Leptongenesis scenario, the more familiar one for the baryon asymmetry is Baryogenesis. The center condition is a first order electroweak phase transition sakharov , is needed to sufficiently maintain a thermal imbalance at a critical temperature during the mass generation of elementary particles. This condition down to calculating the high-temperature effective Higgs potential and investigating the electroweak phase transition (EWPT). The current calculations mostly stop at the one loop level.
Two scenarios are commonly seen when analyzing the structure of EWPT process: the single-stage scenario as in Refs. zb ; pkll ; yukawa1 ; yukawa2 , most theories under this scenario are models that modify SM but do not change the group structure; and mixed or separated multi-phase scenarios plv ; phonglongvan2 ; ptl ; pa , via models have group structures different from SM.
Currently there are three methods to survey the effective potential of EWPT process. The first one in Refs. thdm ; Fuyutob : The effective potential $V_{eff}(v)$ which is a function of $v$ (VEV), written jointly for two VEVs and $v^{2}=v^{2}_{1}+v_{2}^{2}$. The second one in Refs. dori ; mayumi : the effective potential is a function of two VEVs but EWPT is still one phase as SM. The third method is ours, in Ref.pal : the effective potential into two separate parts $[V_{eff}=V_{eff}(v_{1})+V_{eff}(v_{2})]$ and examining two discrete stages.
The two Higgs model with $S_{3}$ symmetry (2HDM-$S_{3}$) 1601 , an extended version of 2HDM, has been studied by us in terms of strigers for a first order electroweak phase transitions pal . Other roles of $S_{3}$ symmetry in the quark sector or the Yukawa couplings, are discussed in Refs. bpal1 ; bpal2 ; bpal3 ; braconi .
The above center condition was solved in the two-Higgs-doublet model with $S_{3}$ symmetry. When adding the $S_{3}$ symmetry to 2HDM, it makes the path of EWPT process discrete. This decoherence is expressed by two discretization stages corresponding to two VEVs. Typically, the first phase goes with $v_{2}$, the second phase goes with $v_{1}$.
2HDM-$S_{3}$
$\Downarrow\text{breaking $v_{2}$}$
$\Downarrow\text{breaking $v_{1}$}$
$U(1)_{Q}$
The $S_{3}$ symmetry acts as the one to simplify the parameters of Higgs potential and can make the EWPT process discrete. But it does not play the dominant role in terms of particle interaction dynamics in 2HDM. Therefore, the value domains of $\tan\beta$ drawn from different data in 2HDM are still valid in 2HDM-$S_{3}$. These remarks are like the following conclusions:
  • According to Ref. beta1 ; beta2 ; bpal1 ; bpal2 ; dori , when calculating decay channels and combining LHC data, $171<\tan\beta<$ 17. Cases of $\tan\beta$ less than one do not exist. The masses of heavy particles are more than 190 GeVmayumi ; CMS2 ; 126 .
  • On the other hand, the independent gauge of EWPT (specifically the effective potential and the phase transition strength) zb ; 1101.4665 ; 1101.4665b ; 1101.4665c ; Arefe . The daisy loops are not the main cause of EWPT but reduces the strength of EWPT r23 . Therefore, when investigating around the electroweak phase transition temperature region, for simplicity we can ignore their contributions.
  • In Refs. 1305.6610 ; 1504.05949 ; PRL ; 125 , the strong first order EWPT in 2HDM with the condition, the hierarchy of mass, $m_{H^{\pm}}<m_{H}<m_{A}$ and $101<\tan\beta<$ 10.
  • The results in Refs. 2111.13079J. High Energy Phys ; 128 confirm that the masses of additional Higgs bosons are typically $300-400$ GeV for a first order EWPT.
  • Furthermore, in Ref. dori ; 124 ; 127 there are many scenarios among the additional Higgs bosons were studied in the EWPT problem, specifically $m_{H^{\pm}}=m_{A}$ or $m_{H}=m_{H^{\pm}}$.
The above conclusions lead to an instruction for surveying 2HDM-$S_{3}$ also with a parameter region, $171\sim\tan\beta<$ 17 and the masses of additional bosons must be larger than 200 GeV. Our result in Ref.pal agrees with the above conclusions and the results in Ref. dori ; davidson ; 129 ; 130 , which concludes that $\tan\beta$ is not a meaningful parameter in 2HDM. This is also the motivation for us to fully study the Baryogenesis scenario in 2HDM-$S_{3}$.
The main results of Ref.pal are outlined as follows. The ratio parameter $\tan\beta$ does not cause the dynamics of phase transition and the effective potential should be written as two components, each corresponding to a VeV. However, $\tan\beta$ determines the value range of addinational particles. The $S_{3}$ symmetry acts as a ”boundary” between the two electroweak phase transitions.
To complete the Baryogenesis scenario in 2HDM-$S_{3}$ and continue the results of Ref.pal , this paper presents the calculations of sphaleron and gravitational wave. Sphaleron is a process associated with the $B$ violation, so calculating and testing its rules to more firmly confirm the solution of baryon asymmetry problem in this model. Gravitational waves (GW) produced in the EWPT are almost a accompanying problem that can be connected to experimental data of current detectors.
The sphaleron energy functional is currently calculated in two different ways (the ansatz with scale-free parameters 10 and the smooth one 11 ; 46 ) but both require numerical solutions, which we summarize in Ref.pkll . Similarly, the key parameters and calculation of gravitational wave energy density are also summarized in Ref.ptl .
In addition, the effect of $S_{3}$ symmetry should be further evaluated even though dynamically the one has no impact. The difference among sphaleron energies and gravitational wave energy densities at the two phase transitions are considered. These bandgaps are sketched as phenomenologically the effects of $S_{3}$ symmetry.
The paper has the following structure. Except for the Introduction (Sec. I) and the Conclusion and Outlooks (Sec. IV), Appendix A and Sec. II discusses the control of first order electroweak phase transition problem. This is a summary of the results in Ref.pal and an analysis of the phase transition strength in values of $a$ to further confirm the independence of $S$ according to $a$. More importantly, we show how to control an first order electroweak phase transition. In Sec. III we derive the sphaleron energy functional and the gravitational wave energy density which undergoes important parameters. Then we solve them and compare the results with existing data or evaluate them.
II Controling EWPT in The 2HDM with $S_{3}$ symmetry
The search for a first order electroweak phase transition is essential, it outlines a possible path to the matter-antimatter asymmetry problem. This has been done and evaluated in Ref.pal . In this section, the conditions for a first order EWPT will be comprehensively evaluated.
II.1 The Higgs potential
The generic scalar potential of 2HDM-$S_{3}$ 1601 ; melic in the complex representation, has a simpler form that contains 5 real parameters:
$V_{\text{Higgs}}^{\text{2HDM}\otimes S_{3}}\equiv\,V(\phi_{1},\phi_{2})=$ $\mu_{1}^{2}(\phi_{2}\phi_{2}^{\dagger}+\phi_{1}\phi_{1}^{\dagger})+\frac{1}{2}l_{1}(\phi_{2}\phi_{2}^{\dagger}+\phi_{1}\phi_{1}^{\dagger})^{2}+\frac{1}{2}l_{2}(\phi_{2}\phi_{2}^{\dagger}-\phi_{1}\phi_{1}^{\dagger})^{2}$
$\qquad+l_{3}(\phi_{1}^{\dagger}.\phi_{2})(\phi_{2}^{\dagger}.\phi_{1})-\mu_{2}^{2}\left(\phi_{1}^{\dagger}.\phi_{2}+\phi_{2}^{\dagger}.\phi_{1}\right)+H.C.$ (1)
The above potential is obtained from adding $S_{3}$ symmetry and the $S_{3}$ soft violating potential to the initial position of 2HDM. The process of calculation and reduction to the above potential is the resetting of scalar fields and parameters presented in Refs.1601 ; melic . Eq. (1) has mixed components between $\phi_{1}$ and $\phi_{2}$ but is much simpler than the original 2HDM potential.
The fields $\phi_{1}$ and $\phi_{2}$ have the following VEVs:
$\langle\phi_{1}\rangle_{0}=\frac{1}{\sqrt{2}}\begin{pmatrix}0\\ v_{1}\end{pmatrix},\langle\phi_{2}\rangle_{0}=\frac{1}{\sqrt{2}}\begin{pmatrix}0\\ v_{2}\end{pmatrix}.$ (2)
For a model with more than one VEV, there is usually a scaling parameter among the VEVs that are included to investigate phenomenological problems. In the model of interest, a very familiar parameter is introduced manually,
$\tan\beta=\frac{s_{\beta}}{c_{\beta}}=\dfrac{v_{2}}{v_{1}}\,,$ (3)
where $s_{\beta}\equiv\sin\beta,\,c_{\beta}\equiv\cos\beta$. $\beta$ is the mixing angle between the charged Higgs and pseudoscalar particle. The value of $\tan\beta$ is summarized in the introduction secction.
When calculating the EWPT problem, the Higgs and gauge boson sector have the strongest contributions. After diagonalizing Eq.(1) and the gauge field sector pal , the masses of Higgs fields are like Eq.(4).
$\begin{split}&m_{H^{\pm}}^{2}=-l_{2}v^{2},\\ &m_{A}^{2}=-\dfrac{1}{2}(2l_{2}-l_{3})v^{2},\\ &m_{H}^{2}=\dfrac{v^{2}}{4}\bigg{[}2l_{1}+l_{3}+\sqrt{16l_{1}l_{2}(c_{\beta}^{2}-s_{\beta}^{2})^{2}-8l_{2}l_{3}+l_{3}^{2}-4l_{1}l_{3}(c_{\beta}^{4}-6c_{\beta}^{2}s_{\beta}^{2}+s_{\beta}^{4})}\bigg{]}=\frac{f_{H}}{4}.v^{2},\\ &m_{h}^{2}=\dfrac{v^{2}}{4}\bigg{[}2l_{1}+l_{3}-\sqrt{16l_{1}l_{2}(c_{\beta}^{2}-s_{\beta}^{2})^{2}-8l_{2}l_{3}+l_{3}^{2}-4l_{1}l_{3}(c_{\beta}^{4}-6c_{\beta}^{2}s_{\beta}^{2}+s_{\beta}^{4})}\bigg{]}=\frac{f_{h}}{4}.v^{2}\,.\end{split}$ (4)
The masses of $H^{\pm}$ and $A$ have no mixing between the two VEVs, this is not possible in the 2HDM model. Similarly, the gauge particles also do not have this mixing. However, we cannot eliminate this mixing for $H$ and $h$ either, so we are forced to approximate $f_{H}$ and $f_{h}$. The mass spectra of them can be derived by expanding the kinetic energy component of the higgs field. The mass spectra of all particles that contribute significantly to the EWPT process in Table 1.
Particles $m(v_{0})$ [GeV] $m^{2}(v_{1},v_{2})$ $m^{2}(v_{1})$ $m^{2}(v_{2})$ $n$
$W^{\pm}$ $80.442$ $\frac{g^{2}v^{2}}{4}$ $\frac{g^{2}v_{1}^{2}}{4}$ $\frac{g^{2}v_{2}^{2}}{4}$ $6$
$Z$ $91.18$ $(g^{2}+g^{\prime 2})\frac{v^{2}}{4}$ $(g^{2}+g^{\prime 2})\frac{v_{1}^{2}}{4}$ $(g^{2}+g^{\prime 2})\frac{v_{2}^{2}}{4}$ $3$
$h$ $125$ $\frac{1}{4}f_{h}v^{2}$ $\frac{1}{4}f_{h}v_{1}^{2}$ $\frac{1}{4}f_{h}v_{2}^{2}$ $1$
$H$- $\frac{1}{4}f_{H}v^{2}$ $\frac{1}{4}f_{H}v_{1}^{2}$ $\frac{1}{4}f_{H}v_{2}^{2}$ $1$
$A$- $-\dfrac{1}{2}(2l_{2}-l_{3})v^{2}$ $-\dfrac{1}{2}(2l_{2}-l_{3})v_{1}^{2}$ $-\dfrac{1}{2}(2l_{2}-l_{3})v_{2}^{2}$ $1$
$H^{\pm}$- $-l_{2}v^{2}$ $-l_{2}v_{1}^{2}$ $-l_{2}v_{2}^{2}$ $2$
$t$ $173.1$ $f_{t}^{2}v^{2}$ $f_{t}^{2}v_{1}^{2}$ $f_{t}^{2}v_{2}^{2}$ $-12$
Table 1: Squared mass of the gauge bosons and scalar bosons in 2HDM-$S_{3}$; whereas mass of the $W^{\pm}$, $Z$ and $t$ is the same as the one in SM; $v^{2}=v_{1}^{2}+v_{2}^{2}$; $v_{0}=246$ GeV.
In this model $v^{2}=v^{2}_{2}+v^{2}_{1}$, we put $v^{2}=a.v^{2}_{2}$. Therefore, the masses of particles in terms of $a$ are given in Table 2.
Particles $m^{2}(v_{1},v_{2})$ $m^{2}(v_{2})$ $m^{2}(v_{1})$
$m^{2}_{W^{\pm}}$ $\frac{g^{2}v^{2}}{4}$ $m^{2}_{W^{\pm}}/a$ $m^{2}_{W^{\pm}}(v_{2}).(a-1)$
$m^{2}_{Z}$ $(g^{2}+g^{\prime 2})\frac{v^{2}}{4}$ $m^{2}_{Z}/a$ $m^{2}_{Z}(v_{2}).(a-1)$
$m^{2}_{h}$ $\frac{1}{4}f_{h}v^{2}$ $m^{2}_{h}/a$ $m^{2}_{h}(v_{2}).(a-1)$
$m^{2}_{H}$ $\frac{1}{4}f_{H}v^{2}$ $m^{2}_{H}/a$ $m^{2}_{H}(v_{2}).(a-1)$
$m^{2}_{A}$ $-\dfrac{1}{2}(2l_{2}-l_{3})v^{2}$ $m^{2}_{A}/a$ $m^{2}_{A}(v_{2}).(a-1)$
$m^{2}_{H^{\pm}}$ $-l_{2}v^{2}$ $m^{2}_{H^{\pm}}/a$ $m^{2}_{H^{\pm}}(v_{2}).(a-1)$
$m_{t}^{2}$ $f_{t}^{2}v^{2}$ $m^{2}_{t}/a$ $m^{2}_{t}(v_{2}).(a-1)$
Table 2: Squared mass of the gauge and scalar bosons in 2HDM-$S_{3}$.
II.2 The temperature effective Higgs potential and cotronling EWPT
With $v_{1}$ different from $v_{2}$, this model has two stages of phase transition and assume $v_{1}<v_{2}$, which means $21<a<$ 2 or $1<\tan\beta$. As the above sections pointed out, in 2HDM-$S_{3}$, the masses of particles can be changed such that there are no mixing terms between $v_{1}$ and $v_{2}$. So the correct effective potential for this model is
$V_{eff}^{S_{3}}=V_{eff}^{S_{3}}(v_{1})+V_{eff}^{S_{3}}(v_{2}).$ (5)
With the above formula for the potential, the phase transition strength does not depend on $a$. This has been proven in Ref.pal . The high-temperature effective potential when considering daisy loops is rewritten as follows,
$V_{eff}(\phi,T)$ $=$ $\lambda(T)\phi^{4}+D(T^{2}-T_{0}^{2})\phi^{2}-\frac{T}{12\pi}\sum_{h,W,Z,H,H^{\pm},A}g_{i}\left[\frac{m_{i}^{2}\phi^{2}}{v^{2}_{b0}}+\Pi_{i}(T)\right]^{\frac{3}{2}}.$ (6)
where, $b=1,2$ then $v_{10},v_{20}$ are $v_{1},v_{2}$ at $T=0$ respectively and
$\lambda(T)$ $=$ $\ \frac{m_{h}^{2}+m_{H}^{2}}{8v^{2}_{b0}}+\frac{1}{64\pi^{2}v^{4}_{b0}}\sum_{h,W,Z,H,H^{\pm},A,t}g_{i}m_{i}^{4}\ln{\left(\frac{A_{i}T^{2}}{m_{i}^{2}}\right)},$ (7)
$D$ $=$ $\ \frac{m_{h}^{2}+6m_{W}^{2}+3m_{Z}^{2}+m_{H}^{2}+2m_{H^{\pm}}^{2}+m_{A}^{2}+6m_{t}^{2}}{24v^{2}_{b0}},$
$T_{0}^{2}$ $=$ $\ \frac{1}{D}\left[\frac{m_{h}^{2}+m_{H}^{2}}{4}-\sum_{h,W,Z,H,H^{\pm},A,t}\frac{g_{i}m_{i}^{4}}{32\pi^{2}v^{2}_{b0}}\right],$
$\Pi_{W}(T)$ $=$ $\ \frac{22}{3}\frac{m_{W}^{2}}{v^{2}_{b0}}T^{2},\quad\Pi_{Z}(T)=\frac{22}{3}\frac{(m_{Z}^{2}-m_{W}^{2})}{v^{2}_{b0}}T^{2},$
$\Pi_{h}(T)$ $=$ $\ \frac{m_{h}^{2}+2m_{W}^{2}+m_{Z}^{2}+2m_{t}^{2}}{4v^{2}_{b0}}T^{2},$
$\Pi_{A}(T)$ $\sim$ $\frac{m_{A}^{2}}{v^{2}_{b0}}T^{2},\quad\Pi_{H^{\pm}}(T)\sim\frac{m_{H^{\pm}}^{2}}{v^{2}_{b0}}T^{2},\quad\Pi_{H}(T)\sim\frac{m_{H}^{2}}{v^{2}_{b0}}T^{2}.$
In Eq.7, $\ln A_{i}=3.907$ or $\ln A_{i}=1.135$ for bosons or fermions. $\Pi_{W}(T)$, $\Pi_{Z}(T)$, $\Pi_{A}(T)$, $\Pi_{H^{\pm}}(T)$, $\Pi_{H}(T)$ are functions that characterize the daisy loop contributions. The daisy loops of charged Higgs $(H^{\pm})$, neutral Higgs $(H)$ and $A$ are omitted because the transition temperatures less than 200 GeV, the contribution of these loops is negligible. This has been clearly demonstrated in the articles pal ; phong2022 .
From many conclusions, we propose a way to investigate or control the Higgs potential to have a first order phase transition or $S=\frac{v_{c}}{T_{c}}>1$ ($v_{c}$ and $T_{c}$ are VEV and critical temperature respectively),
  • Build high temperature effective potential. Select the investigation variable, usually the masses of new particles.
  • Survey in the absence of daisy loops. Specify the range of values for variables to have $S>1$.
  • If the masses of new particles are greater than 200 GeV, there are no need to use additional daisy loops.
Figure 1: The strength of EWPT is a function of $a$. The masses of particles are in GeV. $m_{H^{\pm}},m_{A}$ and $m_{H}$ in the figure are just $m_{H^{\pm}}(v_{2}),m_{A}(v_{2})$ and $m_{H}(v_{2})$.
In Fig. 1, the strength $S$ almost does not change with $a$, for different mass cases. This is another way to confirm the results of previous article (in Ref.pal ). Also in Fig.1, the lines are not absolutely parallel to the $a$ axis (the slopes are very small). This means that $S$ still depends on $a$, because the $\Pi$ functions which are the contributions of daisy loops, still depend on $a$ but are negligible. Through such analysis, also following many conclusions from other authors, it is once again fully confirmed that the contribution of daisy loops in the vicinity of critical temperature $(T_{c}\sim 100-150$ GeV) is negligible.
With different mass pairs, the values of phase transition strength do not vary too much. These masses are arbitrary but falls within the range of values drawn from other studies. Not any mass pair of two particles gives a first order EWPT. The three parameters including $a$ and the two mass parameters are bound together. This is one of the conditions that can be used as a reference for the mass range of new particles. Thus the search for a first order electroweak phase transition depends on the masses of particles and even $a$ (when considering daisy loops).
As in Ref.pal , the value of $a$ is in the range 1 to 2. $a$ (or $\tan\beta$) is the parameter that is entered by hand, in the electroweak phase transition problem, although it has no role in influencing the strength of phase transition, but physically, it characterizes the distance between two phase transitions.
There are two ways to control a first order phase transition: the first, choose the masses of the particles (which must also be large enough), then control $a$ to find $S>1$ as in this article (although the dependence on $a$ is quite small when taking into account daisy loops); the second, which is the way done in Ref. pal . For models with multiple phase transitions, we should do both of these to cross-check the results.
According to Fig.2, the two potential functions corresponding to $v_{2}$ and $v_{1}$ get closer to each other as $a$ increases. With the uniform convergence of second minima (i.e. always less than $246$ GeV), the two effective potentials are always similar to each other. Although Fig.2 only corresponds to one mass pair, it is quite general, other mass pairs are presented in Appendix A.
Replacing the investigation in terms of $\tan\beta$ with $a$ actually gives an advantage in sketching the evolution of two components in the effective potential. Mathematically, the value domain of $\tan\beta$ is an open one (i.e. there is no upper bound). But when examined in terms of $a$, the value domain of $a$ is closed, it lies between 1 and 2. This change is like a mapping of $\tan\beta$ into $a$. In summary, finding a first order electroweak phase transition with an effective potential pair in 2HDM-$S_{3}$ is mathematically and physically feasible.
Figure 2: The effective potential $V_{eff}$ at $T=T_{c}$. $m_{H^{\pm}}(v_{2})=m_{A}(v_{2})=350$ GeV, $m_{H}(v_{2})=160$ GeV. The red line corresponds to $V_{eff}(v_{2})$. The orange line corresponds to $V_{eff}(v_{1})$.
In Fig.2, the non-zero minimum of effective potential corresponding to $v_{2}$ is always larger than that of $v_{1}$. The yellow lines in Fig.2 approach the red lines as $a$ increases or the two phase transitions get closer together. When $a=2$, the two potential functions coincide, but then the electroweak phase transition is the simultaneous phase transition of two VEVs. This corresponds to the existence of two scalar Higgs fields in this model. It is interesting to note that these two scalar fields are partly involved in both phase transitions. No one scalar particle is involved in only one phase transition. The nature of this problem is that the higgs potential contains up to two Higgs doublets and is a general mixture of these two doublets. Furthermore, in Table 3, for a given value of $a$, $S$ of the two EWPT stages are the same. This confirms the results in Ref.pal . However, to find the maximum value of $S$ which is often not very precise, the method in Ref.pal gives better results.
a $v_{c}$ [GeV] $T_{c}$ [GeV] $S$ $E$ [GeV] $\alpha$ $\frac{\beta}{H*}$ $\Omega h^{2}(f_{peak})\times 10^{-14}$
1.1 $v_{2}$=154.618 159.443 0.969736 7841.47 0.017436 22.1311 1.82991
$v_{1}$=48.8945 50.4204 0.969735 2479.69 0.0174359 22.1311 1.8299
1.2 $v_{2}$=153.951 152.784 1.00764 7588.49 0.019334 22.3506 2.70965
$v_{1}$=68.8489 68.3271 1.00764 3393.67 0.019334 22.3506 2.70965
1.3 $v_{2}$=153.12 146.913 1.04225 7365.48 0.0211846 22.5607 3.82934
$v_{1}$=83.8672 80.4677 1.04225 4034.24 0.0211846 22.5607 3.82934
1.4 $v_{2}$=152.188 141.682 1.07415 7166.8 0.02299 22.7626 5.21179
$v_{1}$=96.2521 89.6077 1.07415 4532.68 0.02299 22.7626 5.21179
1.5 $v_{2}$=151.2 136.979 1.10382 6988.23 0.0247537 22.9576 6.87871
$v_{1}$=106.914 96.8587 1.10382 4941.43 0.0247537 22.9576 6.87871
1.6 $v_{2}$=150.187 132.717 1.13164 6826.53 0.0264803 23.1465 8.85171
$v_{1}$=116.335 102.802 1.13164 5287.81 0.0264803 23.1465 8.85171
1.7 $v_{2}$=149.173 128.83 1.15791 6679.16 0.0281746 23.3302 11.1534
$v_{1}$=124.807 107.787 1.15791 5588.19 0.0281746 23.3302 11.1534
1.8 $v_{2}$=148.175 125.264 1.1829 6544.11 0.0298422 23.5091 13.8084
$v_{1}$=132.531 112.039 1.1829 5853.23 0.0298422 23.5091 13.8084
1.9 $v_{2}$=147.203 121.977 1.20681 6419.72 0.0314887 23.6838 16.8449
$v_{1}$=139.649 115.717 1.20681 6090.28 0.0314887 23.6838 16.8449
2 $v_{2}$=146.268 118.934 1.22982 6304.68 0.03312 23.8544 20.2955
$v_{1}$=146.268 118.934 1.22982 6304.68 0.03312 23.8544 20.2955
Table 3: Results in case $m_{H^{\pm}}(v_{2})=m_{A}(v_{2})=350$ GeV, $m_{H}(v_{2})=160$ GeV.
III Sphalerons and Gravitational waves in the 2HDM-$S_{3}$
The 2HDM model with the $S_{3}$ symmetry shows a ”twin bubble nucleation or double bubble nucleation”. The first possible demonstration of this is that the masses of particles depends on the two VEVs. Two bubbles can form at the same time or one after the other. The following numerical analysis will clarify that.
III.1 Sphaleron
The sphaleron energy functional consists of three components: the contribution of gauge fields, Higgs kinetic energy and effective potential.
[…]
The contributions of exotic Higgs bosons ($H^{\pm}$, $A$ and $H$) can be ignored. However, they still contribute to the effective potential components. The static field approximation (ie., $W^{a}_{0}=0$) has been assumed and sphaleron has a spherically symmetric form 11 :
$\left\{\begin{array}[]{ll}\phi(r)=\frac{v_{0}}{\sqrt{2}}h(r)in_{a}\sigma\left(\begin{array}[]{cc}0\\ 1\end{array}\right),\\ W_{i}^{a}(r)=\frac{2}{g}\epsilon^{aij}n_{j}\frac{f(r)}{r},\end{array}\right.$ (9)
where $n_{i}\equiv\frac{x^{i}}{r}$ and $r$ is the radial coordinate in the spherical coordinates. The functions $h(r),f(r)$ are profile functions, which characterize bubbles.
The Sphaleron energy functional at temperature $T$ is written as follows
[…]
where $g^{2}=\frac{G_{F}8m^{2}_{W}}{\sqrt{2}}$; $G_{F}=1.166\times 10^{-}{15}\,\textrm{GeV}^{-2}$; $m_{W}=80.39\,\textrm{GeV}$, $\xi\equiv gv_{0}r$.
To make the above function converge in the process of solving, the last term is rewritten as a difference. Thus specifically for each stage will be
[…]
$+\frac{\xi^{2}}{g^{2}v_{b0}^{4}}(V_{eff}(v_{c,b}h_{b},T_{c,b})-V_{eff}(v_{c,b},T_{c,b}))\bigg{]}.$ (11)
Taking the variation of above function in terms of $h$ and $f$, after minimizing the Sphaleron functional, the equations of motion are obtained,
$\frac{d^{2}f_{b}}{d\xi^{2}}$ $=\frac{2}{\xi^{2}}f_{b}(1-f_{b})(1-2f_{b})-\frac{1}{4}(1-f_{b})h_{b}^{2},$ (12)
$\frac{d^{2}h_{b}}{d\xi^{2}}$ $=\frac{-2}{\xi}\frac{dh_{b}}{d\xi}+\frac{2}{\xi^{2}}h_{b}(1-f_{b})^{2}+\frac{1}{g^{2}v_{b0}^{4}}\frac{\partial V_{eff}(v_{c,b}h_{b},T_{c,b})}{\partial h_{b}}.$ (13)
Here, $h_{b}(\xi)$ and $f_{b}(\xi)$ have boundary conditions,
$\left\{\begin{array}[]{ll}h_{b}(\xi\to 0)=f_{b}(\xi\to 0)=0,\\ h_{b}(\xi\to\infty)=f_{b}(\xi\to\infty)=1.\end{array}\right.$ (14)
Note that in the above equations, the run index $b=1,2$ is the phase transition of $v_{1,2}$, respectively. From Eqs. (12,13,14), the numerical solutions are obtained as Fig.3.
Figure 3: $m_{H^{\pm}}(v_{2})=m_{A}(v_{2})=350$ GeV, $m_{H}(v_{2})=160$ GeV.
In Fig.3, the blue lines approach the green lines, the red lines go to the orange lines, as $a$ increases. The blue and green lines represent the bubbles of gauge fields, the orange and red lines represent the bubbles of Higgs fields, corresponding to the two phases.
When the phase transition occurs, $v_{2}$ is broken before $v_{1}$. The blue and red lines (corresponding to the bubbles associated with $v_{2}$) approach 1 faster or erlier than the green and orange lines (corresponding to the bubbles associated with $v_{1}$).
In other words, the process of forming bubble nucleations, to give the masses of particles, from massless one: the first they form a bubble nucleation corresponding to $v_{2}$ first, then form a bubble nucleation corresponding to $v_{1}$ so the initial particles form a pair of bubble nucleation. But these two nucleations can form at the same time when $a=2$, we can name this pair as a twin bubble nucleation.
And the sphaleron energy as the fifth column in Table 3. According to Appendix A, $a$ and the sphaleron energy both increases. The total energy value of these two phase transitions is in the range of 10 to 15 TeV.
III.2 Gravitational waves
The formulas to calculate the gravitational wave energy density (GWED) will be systematized as a setting for the calculation process. There are three processes in the production of GWs with a first-order EWPT: the bubble wall collisions (Coll), sound waves (SW) and Magnetohydrodynamic (MHD) turbulence 71 . Usually they are added linearly to calculate the gravitational wave energy density 71 :
$h^{2}\Omega_{GW}=h^{2}\Omega_{Coll}+h^{2}\Omega_{sw}+h^{2}\Omega_{tur}.$ (15)
The GWED of SW depends on $\alpha$ and $\beta/H^{*}$ phu5gw ; 71 :
$h^{2}\Omega_{sw}(f)\simeq 2.65\times 10^{-6}v_{b}k^{2}_{w}\left[\frac{H_{*}}{\beta}\right]\left[\frac{\alpha}{1+\alpha}\right]^{2}\left[\frac{100}{g^{*}}\right]^{\frac{1}{3}}S_{sw}(f),$ (16)
where
$S_{sw}(f)=(f/f_{sw})^{3}\left(\frac{7}{4+3(f/f_{peak.sw})^{2}}\right)^{7/2},$ (17)
and at the temperature $T_{*}$, the nucleation temperature 71 , the peak frequency estimated as follows
$f_{peak.sw}=1.9\times 10^{-2}mHz\times\frac{1}{v_{b}}\frac{\beta}{H^{*}}\frac{T_{*}}{100}\left(\frac{g_{*}}{100}\right)^{1/6}.$ (18)
The bubble wall velocity is defined via $\alpha$ 71 :
$v_{bn}=\frac{\frac{1}{\sqrt{3}}+\sqrt{\alpha^{2}+2\frac{\alpha}{3}}}{1+\alpha}.$ (19)
A good approximation is $v_{bn}\sim 1$ from the value of $\alpha$ in Table 3. Therefore, the generation of gravitational waves has only two components: the sound wave and turbulence 71 ; 73 ; 73a ; 73b ; 73c ; 73d ; 73e .
GWs from turbulence,
$h^{2}\Omega_{tur}(f)\simeq 3.35\times 10^{-4}\left(\frac{H^{*}}{\beta}\right)v_{b}\left(\frac{k_{t}\alpha}{1+\alpha}\right)^{\frac{3}{2}}\left(\frac{100}{g^{*}}\right)^{1/3}S_{tur}(f),$ (20)
in which
$S_{tur}(f)=\frac{(f/f_{peak.tur})^{3}}{\left(1+(f/f_{peak.tur})\right)^{11/3}\left(1+8\pi f/h^{*}\right)},$ (21)
and at $T_{*}$, Eq. (21) reads
$h_{*}$ $=$ $16.5\times 10^{-3}mHz\times\frac{T_{*}}{100}\left(\frac{g_{*}}{100}\right)^{1/6}\,,$
$f_{peak.tur}$ $=$ $2.7\times 10^{-2}mHz\times\frac{1}{v_{b}}\frac{\beta}{H^{*}}\frac{T_{*}}{100}\left(\frac{g_{*}}{100}\right)^{1/6}.$ (22)
One can estimate 71 that
$k_{t}=(0.05-0.1)\times k_{w};\quad k_{w}=\alpha\left[0.73+0.083\sqrt{\alpha}+\alpha\right]^{-1}\quad\text{when}\quad v_{bn}\sim 1.$ (23)
The next, by using a possible approximation phong2022 , because when calculating GW we only consider the effective potential and scalar field component 71 ; phu5gw ,
$\frac{\beta}{H^{*}}\approx\left[T\frac{d\left(0.45\frac{E^{T}_{sph}}{T}\right)}{dT}\right]_{T=T_{C}},$ (24)
in which $\frac{\beta}{H^{*}}$ is a dimensionless quantity and is calculated using the sphaleron energy and the potential energy. Additionally, in Ref. 8gw , at $T_{*}=T_{c}$, $H^{*}$ is estimated as
$H^{*}$ $=$ $106.75\displaystyle 4.5\times 10^{-22}\left[\frac{T_{c}}{MeV}\right]^{2}\,\textrm{ MeV}\left[\frac{g_{*}}{10.75}\right]^{2};\,g^{*}=$ 106.75 . (25)
The last parameter is $\alpha$ 2gw ; 2gwa ; 2gwb , summarized as the below formulas:
$\alpha$ $=\frac{\epsilon}{\rho_{rad}^{*}},\rho_{rad}^{*}=g^{*}\pi^{2}\frac{T^{4}_{c}}{30}=106.75\pi^{2}\frac{T^{4}_{c}}{30},$ (26)
$\epsilon$ $=\left(V_{eff}(v(T),T)-T\frac{d}{dT}V_{eff}(v(T),T)\right)_{T=T_{C}}.$ (27)
$\alpha$ depends almost entirely on the effective potential. Therefore, in models, for a one-loop effective potential, $\alpha$ can be estimated under a value of $S$ and GW can be calculated at $T_{c}$.
Figure 4: Gravitational wave energy density at $T_{*}=T_{C}$. The blue line corresponds to $v_{2}$, the purple line corresponds to $v_{1}$. $m_{H^{\pm}}(v_{2})=m_{A}(v_{2})=600$ GeV, $m_{H}(v_{2})=300$ GeV.
The two lines in each sub-figure in Fig.4 go closer to each other as $a$ increases. This is also consistent with the previous observations about the two EWPT stages.
From the numerical results in Table 3 and Appendix A, the gravitational wave energy density of two phase transitions are equal at $f=f_{peak}$ at $T_{c}$. This is shown visually in Fig. 4. Because $a$ does not affect the phase transition strength of the two stages (i.s. the phase transition strength of the two stages is equal for one value of $a$). $\Omega h^{2}$ ranges widely, from $10^{-14}$ to $10^{-11}$. As $a$ increases, $\Omega h^{2}$ increases sharply.
The gravitational wave detection frequency is solved in the range below a few mHz. At such low frequencies there is currently no experimental data. However, future detectors may be able to detect gravitational energy densities in the small frequency range as summarized in Table 4. The LISA and DECIGO detectors can pick up the gravitational wave signal generated by the EWPT process in the 2HDM-$S_{3}$ model.
$f[mHz]$ $\Omega h^{2}$ Kind Refs.
$0.02-0.12$ $2\times 10^{-12}-10^{-10}$ LISA Refs.phu1gw ; kudoh ; thranel
$0.01-0.12$ $5\times 10^{-15}-8\times 10^{-14}$ DECIGO Refs.phu1gw ; kudoh ; thranel
$0-0.2$ Not yet BBO Refs.phu1gw ; kudoh ; thranel
Table 4: Sensitivity of proposed GW.
IV CONCLUSION AND OUTLOOKS
$a$ $v_{b0}$ [GeV] $v_{c}$ [GeV] $E_{0}$ [GeV] $E$ [GeV] $E/v_{c}$ $E_{b0}/v_{b0}$ Difference
1.3 $v_{20}=215.592$ $v_{2}$=153.12 7902.28 7365.48 48.10 36.65 $24\%$
$v_{10}=118.656$ $v_{1}$=83.8672 4302.69 4034.24 48.10 36.26
1.5 $v_{20}=200.96$ $v_{2}$=151.2 7311.87 6988.23 46.22 36.38 $21\%$
$v_{10}=142.026$ $v_{1}$=106.914 5172.53 4941.43 46.22 36.42
1.7 $v_{20}=188.665$ $v_{2}$=149.173 6826.74 6679.16 44.77 36.18 $19\%$
$v_{10}=157.979$ $v_{1}$=124.807 5707.53 5588.19 44.77 36.13
Table 5: $m_{H^{\pm}}(v_{2})=m_{A}(v_{2})=350$ GeV, $m_{H}(v_{2})=160$ GeV.
According to the data tables, at the same value of $a$, the ratio $E_{b}/v_{b}$ is always equal. By the linear scaling law of sphaleron energy scaling1 ; scaling2 ; scaling3 , $E_{b}/v_{b}=E_{b0}/v_{b0}=const$ ($E_{b0}$ is the sphaleron energy at $T=0$, Eq. (13) calculated by replacing the effective potential at temperature $T$ with one at 0), the evolution (increase or decrease of sphaleron energy) in the two phase transitions is the same. However, $a$ and this progression increase proportionally.
Additionally to test the scaling law (however, this was initially only an approximation to solve the Sphaleron energy problem scaling2 ; scaling3 ). We estimate some cases like Table 5, $E_{b}/v_{b}$ and $E_{b0}/v_{b0}$ differ by no more than $24\%$. Thus the results are not very compatible with the scaling law or this law is only true in a non-zero temperature range but is not true in the temperature range from 0 to $T_{c}$. For some other cases the calculation also breaks this law as in Ref.phascaling . Because there are many triggers outside SM and energy scales other than SM. Furthermore, the solutions $h(\xi)$ and $f(\xi)$ in Eqs. (13) will approach $1$ very slowly at 00, leading to $E_{0}$ no longer holding the Scaling law. Furthermore, since $v_{eff}(T)$ is only used for non-zero temperatures, that is, $V_{eff}(0)\neq\lim_{T\longrightarrow 0}v_{eff}(T)$.
The total sphaleron energy of two phase transitions is about $10-13$ TeV in this model. The value differs by about $1-5$ TeV from the SM. We also need to check the decoupling condition of sphaleron energy decoupling ; decouplingb ; decouplingc as the following formula,
$\frac{E_{sph}(T)}{T}-7ln\left(\frac{v(T)}{T}\right)+ln\left(\frac{T}{100GeV}\right)>(35.9-42.8).$ (28)
Based on the data tables in Appendix A, the $E/T$ ratio is always greater than $50$; $v_{c}$ is not much bigger than $T_{c}$ and almost $T_{c}$ are larger than $100$ GeV so $[-7\ln(v_{c}/T_{c})+\ln(T/100)]$ has a negative value but not less than $-5$. Thus the sphaleron energy always satisfies the decoupling condition in Eq. (28). Also, the nucleation temperature $T_{N}$ is not much larger than $T_{C}$, the supercooling process is small Fuyuto ; Fuyutob ; Fuyutoc . In other words, the onset of phase transition is not so slow. However, the smaller $a$ is, the larger the gap between two phases is, the longer the EWPT is, another supercooling process in this model.
$\beta/H^{*}$ is actually quite difficult to calculate, it is estimated by us from the different contributions to the sphaleron energy. Based on numerical calculations, the contribution of gauge field component to this energy is $55\%$ so having an approximation as Eq.(24). In general an approximation, $\frac{\beta}{H^{*}}\approx T\frac{d}{dT}\left(\gamma.\frac{E^{T}_{sph}}{T}\right)$ can be choosed for other models; $\gamma$ depends on the contributions of components.
In the future, with new experimental data from the LHC, simply confirming the signals of particles beyond SM or sphaleron, it will indirectly confirm the Baryogenesis scenario.
Thus the baryogenesis scenario in the 2HDM-$S_{3}$ model is fully calculated. One assertion is that this model has enough triggers to generate a first order electroweak phase transition, with sphaleron energies around the mid-range value of about 10 TeV which are checked by theoretical conditions and GW energies are compared with experimental possibilities. Additionally, the prospects for GW, which can be used for further reference, found in Ref.gwhite .
Our future work follows the results, calculating CP violation, searching and connecting the strength of EWPT, sphalerons or gravitational waves with experimental data.
ACKNOWLEDGMENTS
This research is funded by Vietnam National Foundation for Science and Technology Development (NAFOSTED) under grant number 103.01-2023.16.
Appendix A Table of data
a $v_{c}$ [GeV] $T_{c}$ [GeV] $S$ $E$ [GeV] $\alpha$ $\frac{\beta}{H*}$ $\Omega h^{2}(f_{peak})\times 10^{-15}$
1.1 $v_{2}$=128.327 166.643 0.770071 7621.73 0.0101885 20.5816 2.40066
$v_{1}$=40.5805 52.697 0.770071 2410.2 0.0101885 20.5816 2.40066
1.2 $v_{2}$=128.655 159.876 0.804718 7346.95 0.0114752 20.6793 3.81175
$v_{1}$=57.5363 71.4987 0.804717 3285.66 0.0114752 20.6793 3.81173
1.3 $v_{2}$=128.713 153.922 0.836219 7109.38 0.0127468 20.7847 5.727
$v_{1}$=70.4988 84.3067 0.836219 3893.97 0.0127468 20.7847 5.72701
1.4 $v_{2}$=128.568 148.628 0.86503 6901.35 0.0139997 20.8951 8.22445
$v_{1}$=81.3135 94.0008 0.86503 4364.8 0.0139997 20.8951 8.22446
1.5 $v_{2}$=128.272 143.878 0.891532 6717.06 0.0152319 21.0086 11.3772
$v_{1}$=90.7021 101.737 0.891532 4749.68 0.0152318 21.0086 11.3772
1.6 $v_{2}$=127.864 139.583 0.916043 6552.06 0.016442 21.1231 15.2531
$v_{1}$=99.043 108.12 0.916043 5075.2 0.016442 21.1231 15.2531
1.7 $v_{2}$=127.373 135.672 0.93883 6402.96 0.0176299 21.2375 19.9147
$v_{1}$=106.568 113.511 0.93883 5357.1 0.0176299 21.2375 19.9147
1.8 $v_{2}$=126.821 132.089 0.960119 6267.14 0.0187957 21.3509 25.4201
$v_{1}$=113.432 118.144 0.960119 5605.5 0.0187957 21.3509 25.4201
1.9 $v_{2}$=126.226 128.789 0.980101 6142.58 0.01994 21.4628 31.8231
$v_{1}$=119.748 122.18 0.980101 5827.37 0.01994 21.4628 31.8231
2 $v_{2}$=125.602 125.735 0.998942 6027.68 0.0210635 21.5728 39.1746
$v_{1}$=125.602 125.735 0.998942 6027.68 0.0210635 21.5728 39.1746
Table 6: Results in case $m_{H^{\pm}}(v_{2})=m_{A}(v_{2})=300$ GeV, $m_{H}(v_{2})=150$ GeV.
a $v_{c}$ [GeV] $T_{c}$ [GeV] $S$ $E$ [GeV] $\alpha$ $\frac{\beta}{H*}$ $\Omega h^{2}(f_{peak})\times 10^{-13}$
1.1 $v_{2}$=198.272 171.699 1.15477 8634.25 0.0319842 22.6293 1.87205
$v_{1}$=62.6992 54.2959 1.15477 2730.39 0.0319842 22.6293 1.87205
1.2 $v_{2}$=195.78 164.759 1.18829 8377.36 0.0346998 22.8808 2.52906
$v_{1}$=87.5557 73.6823 1.18829 3746.47 0.0346998 22.8808 2.52906
1.3 $v_{2}$=193.537 158.617 1.22015 8150.04 0.037412 23.1218 3.33497
$v_{1}$=106.004 86.8783 1.22015 4463.96 0.037412 23.1218 3.33497
1.4 $v_{2}$=191.544 153.149 1.25071 7947.12 0.0401542 23.3511 4.32144
$v_{1}$=121.143 96.86 1.25071 5026.2 0.0401542 23.3511 4.32144
1.5 $v_{2}$=189.799 148.262 1.28016 7764.6 0.0429607 23.5669 5.53165
$v_{1}$=134.208 104.837 1.28016 5490.4 0.0429607 23.5669 5.53165
1.6 $v_{2}$=188.289 143.888 1.30859 7599.4 0.0458661 23.7667 7.02374
$v_{1}$=145.848 111.455 1.30859 5886.47 0.0458661 23.7667 7.02374
1.7 $v_{2}$=186.999 139.975 1.33595 7449.03 0.0489043 23.9476 8.87484
$v_{1}$=156.455 117.111 1.33595 6232.31 0.0489043 23.9476 8.87484
1.8 $v_{2}$=185.909 136.483 1.36213 7311.46 0.0521068 24.1066 11.1852
$v_{1}$=166.282 122.074 1.36213 6539.57 0.0521068 24.1066 11.1852
1.9 $v_{2}$=184.994 133.381 1.38696 7184.98 0.0555005 24.2407 14.0822
$v_{1}$=175.5 126.536 1.38696 6816.27 0.0555005 24.2407 14.0822
2 $v_{2}$=184.227 130.638 1.41021 7068.15 0.0591052 24.3471 17.7223
$v_{1}$=184.227 130.638 1.41021 7068.15 0.0591052 24.3471 17.7223
Table 7: Results in case $m_{H^{\pm}}(v_{2})=m_{A}(v_{2})=500$ GeV, $m_{H}(v_{2})=250$ GeV.
a $v_{c}$ [GeV] $T_{c}$ [GeV] $S$ $E$ [GeV] $\alpha$ $\frac{\beta}{H*}$ $\Omega h^{2}(f_{peak})\times 10^{-12}$
1.1 $v_{2}$=226.219 173.466 1.30411 9149.73 0.0453932 23.7359 0.676314
$v_{1}$=71.5367 54.8549 1.30411 2893.4 0.0453932 23.7359 0.676314
1.2 $v_{2}$=224.01 166.748 1.34341 8891.16 0.0497831 23.9945 0.94705
$v_{1}$=100.18 74.5718 1.34341 3976.25 0.0497831 23.9945 0.94705
1.3 $v_{2}$=222.28 161.054 1.38016 8663.24 0.0545297 24.2059 1.32034
$v_{1}$=121.748 88.213 1.38016 4745.05 0.0545297 24.2059 1.32034
1.4 $v_{2}$=220.939 156.277 1.41376 8460.32 0.0597024 24.3614 1.83858
$v_{1}$=139.734 98.8385 1.41376 5350.77 0.0597024 24.3614 1.83858
1.5 $v_{2}$=219.885 152.323 1.44354 8277.83 0.0653321 24.4547 2.55674
$v_{1}$=155.482 107.709 1.44354 5853.31 0.0653321 24.4547 2.55674
1.6 $v_{2}$=219.015 149.102 1.4689 8112.05 0.0713998 24.4827 3.54026
$v_{1}$=169.648 115.494 1.4689 6283.56 0.0713998 24.4827 3.54026
1.7 $v_{2}$=218.235 146.522 1.48943 7959.9 0.0778348 24.4464 4.85909
$v_{1}$=182.588 122.589 1.48943 6659.73 0.0778348 24.4464 4.85909
1.8 $v_{2}$=217.466 144.494 1.50501 7818.9 0.0845259 24.3505 6.57865
$v_{1}$=194.508 129.24 1.50501 6993.43 0.0845259 24.3505 6.57865
1.9 $v_{2}$=216.653 142.93 1.5158 7687.09 0.0913395 24.2019 8.74961
$v_{1}$=205.535 135.596 1.5158 7292.61 0.0913395 24.2019 8.74961
2 $v_{2}$=215.759 141.75 1.52211 7562.95 0.0981397 24.0094 11.3997
$v_{1}$=215.759 141.75 1.52211 7562.95 0.0981397 24.0094 11.3997
Table 8: Results in case $m_{H^{\pm}}(v_{2})=m_{A}(v_{2})=600$ GeV, $m_{H}(v_{2})=300$ GeV.